On July 31, 20x2, Bismarck Company engaged Duval Tooling Company to construct a special-purpose piece of factory machinery. Construction began immediately and was completed on November 1, 20x2. To help finance construction on July 31 Bismarck issued a $400,000 3 year, 12% note payable at Wellington National Bank, on which interest is payable each July 31. $300,000 of the proceeds of the note was paid to Duval on July 31. The remainder of the proceeds was temporarily invested in short-term marketable securities (debt investments) at 10% until November 1. On November 1, Bismarck made a final $100,000 payment to Duval. Other than the note to Wellington, Bismarck's only outstanding liability at December 31, 20x2 is a $30,000, 8%, 6-year note payable, dated January 1, 20x0, on which interest is payable each December 31. Required: 1) Calculate the interest revenue, weighted-average accumulated expenditures, and total interest cost to be capitalized in 20x2. Round all computations to the nearest dollar and show your work. 2) Prepare the journal entries needed on the books of Bismarck Company at each of the following dates: a) July 31, 20x2 b) November 1, 20x2 3) December 31, 20x2
Answer #1

Answer:

itauicd dui-j 2612 and totat Cos to be Capi In-iere, vevenue $250o weinhtt d avevaqe acum Avu i dabl e (nAcregt, 9000 그n-f v
Answer #2

On July 31, 20x2, Bismarck Company engaged Duval Tooling Company to construct a special-purpose piece of factory machinery. C
Answer #3
july 31 300.000 × 3/13 = 75000 nov 1 100.00p × 0=0 weighted average accumulated expenditure = 75000 interest revenue 100.000 × 10%× 3/ 12= 2500 Avoidable interest= interest capitalize= 75000× 12%= 9000 Actual interest 22.400 b. july 31 db.cash 400.000 cr. Note payable 400.000 db. Machinery 300.000 db. Trading securities 100.000 cr. Cash 400.000
answered by: alda